### v4

# Existence of cohomogeneity one Einstein metrics

Abstract.
This paper derives a sufficient condition for the existence of cohomogeneity one Einstein metrics on double disk bundles of two summands type. The condition is an inequality that involves geometric data from the principal orbits.
2020 Mathematics Subject Classification: 53C25 (primary). Keywords: Einstein metric, cohomogeneity one metric. The author is supported by the NSFC (No. 12071489, No. 12301078), the Natural Science Foundation of Jiangsu Province (BK-20220282), and the XJTLU Research Development Funding (RDF-21-02-083).
1. Introduction
A Riemannian metric $g$ is Einstein if $\mathrm{Ric}(g)=\Lambda g$ for some constant $\Lambda$. The search for Einstein metrics continues to be a challenging problem. A milestone in this field is the existence theory of Kähler–Einstein metrics, with fundamental results including the resolution of the Calabi conjecture in the non-positive case [Aub76, Yau78] and the Yau–Tian–Donaldson conjecture for Fano manifolds [CDS15, Tia15]. While these results have profoundly shaped our understanding of Einstein metrics, many challenges remain in the non-Kähler setting. Significant efforts have been devoted to constructing Einstein metrics in the real Riemannian context. One fruitful approach involves considering homogeneous spaces, where the high degree of symmetry simplifies the Einstein equation to an algebraic system. In the case of homogeneous spaces, a general existence theory has been developed in [WZ86, BWZ04, Böh04]. It is known in [BK06] that each compact simply connected homogeneous space of dimension less than $12$ admits a homogeneous Einstein metric. For low dimensional homogeneous Einstein manifolds up to dimension $7$; see [Jen69, Gor77, ADF96, NR99, NR03, Nik04].
To explore Einstein metrics with large symmetry but beyond homogeneity, a natural direction is cohomogeneity one Einstein metrics, where the isometry group acts with principal orbits $\mathsf{G}/\mathsf{K}$ of codimension one. The cohomogeneity one condition transforms the Einstein equation to a system of ODEs. A landmark example is Page’s metric on $\mathbb{CP}^{2}\sharp\overline{\mathbb{CP}}^{2}$, the firstly known inhomogeneous positive Einstein metric [Pag78]. Page’s metric is generalized in a series of works including [BB82, KS86, PP87, KS88, WW98]. In [Böh98], new cohomogeneity one Einstein metrics were found on some low dimensional manifolds, including spheres and $\mathbb{HP}^{2}\sharp\overline{\mathbb{HP}}^{2}$. The latter space is the quaternionic analogue of Page’s metric. Böhm’s metric is generalized to all high dimensional $\mathbb{HP}^{m+1}\sharp\overline{\mathbb{HP}}^{m+1}$ in [Chi24], confirming numerical evidence in [PP86].
This paper further extends the existence result in [Chi24] to a broader class of closed cohomogeneity one manifolds. Specifically, the principal orbit is the total space of a homogeneous fibration
(1.1) $\mathbb{S}^{d_{1}}=\mathsf{H}/\mathsf{K}\hookrightarrow\mathsf{G}/\mathsf{K}\rightarrow\mathsf{G}/\mathsf{H}=\mathcal{Q}^{d_{2}}.$
The compact cohomogeneity one manifold $M$ of dimension $n+1=d_{1}+d_{2}+1$ is an $\mathbb{S}^{d_{1}+1}$-bundle over $\mathcal{Q}^{d_{2}}$. Since the manifold $M$ is formed by gluing two copies of a disk bundle over $\mathcal{Q}^{d_{2}}$ along $\mathsf{G}/\mathsf{K}$, it is also referred to as a double disk bundle associated to the group triple $(\mathsf{K},\mathsf{H},\mathsf{G})$. We consider double disk bundles of two summands type, where the cohomogeneity one Einstein equation admits an ansatz consisting of two metric components $f_{1}^{2}$ and $f_{2}^{2}$, respectively for $\mathfrak{h}/\mathfrak{k}$ and $\mathfrak{g}/\mathfrak{h}$. Cases where the isotropy representations of principle orbits split into two irreducible summands are included. The cohomogeneity one Einstein equation on $M$ is characterized by the structural triple $(d_{1},d_{2},A)$ associated with $\mathsf{G}/\mathsf{K}$, where the constant $A\geq 0$ is the norm square of the fundamental tensor in the theory of Riemannian submersion [O’N66]. With such a geometric setting, we prove the following existence theorem.
Theorem 1.1.
For any $(d_{1},d_{2})$ with $d_{2}\geq d_{1}\geq 2$, there exists a constant $\chi_{d_{1},d_{2}}\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right]$ such that if $\mathsf{G}/\mathsf{K}$ is a principal orbit with the structural triple $(d_{1},d_{2},A)$ and $A\in(0,\chi_{d_{1},d_{2}})$, there exists at least one cohomogeneity one Einstein metric on $M$.
The case where $d_{1}=1$ was well studied in [BB82] and many aforementioned works that followed, where it was not even assumed that the hypersurface of $M$ is a homogeneous space. We hence consider $d_{1}\geq 2$ in this paper. Note the condition $d_{2}\geq d_{1}$ does not restrict generality. If $d_{2}<d_{1}$, then by slice theorem, a tubular neighborhood around $\mathcal{Q}^{d_{2}}$ is identified as a vector bundle whose rank is larger than $d_{2}$. Therefore, the slice representation of $\mathsf{H}$ contains at least two trivial summands, meaning $\mathsf{H}$ does not act transitively on $\mathsf{H}/\mathsf{K}$. By the assumption of $\mathsf{G}/\mathsf{K}$ being of two summands type, the principal orbit is necessarily a Riemann product $\mathbb{S}^{d_{1}}\times\mathcal{Q}^{d_{2}}$ and $A=0$. Then $M=\mathbb{S}^{d_{1}+1}\times\mathcal{Q}^{d_{2}}$ obviously admits a product metric that is Einstein.
The constant $\chi_{d_{1},d_{2}}$ is determined by minimizing certain rational functions. The explicit formula of $\chi_{d_{1},d_{2}}$ for a general $(d_{1},d_{2})$ is fairly complicated. In an upcoming paper [Chi25b], we derive some estimates of $\chi_{d_{1},d_{2}}$ to provide some concrete new examples of compact Einstein manifolds.
This paper is organized as follows. In Section 2, we introduce the dynamical system for cohomogeneity one Einstein metrics of two summands type. In Section 3, we perform local analysis on critical points representing the initial and terminal conditions. In Section 4, we generalize the global analysis in [Chi24] and prove the main theorem. For the sake of readability, formula and computations obtained by Maple are presented in the Appendix.
2. The cohomogeneity one Einstein equation
For a group triple $(\mathsf{K},\mathsf{H},\mathsf{G})$ as in (1.1), let $\mathsf{G}$ be compact and semisimple. Let $\tilde{b}$ a normal homogeneous metric on $\mathsf{G}/\mathsf{K}$. If $\mathsf{G}$ is simple, let $\tilde{b}$ be the standard homogeneous metric induced by the negative Killing form $-B_{\mathfrak{g}}$. Assume that there exists a $\tilde{b}$-orthogonal decomposition $\mathfrak{g}=\mathfrak{h}/\mathfrak{k}\oplus\mathfrak{g}/\mathfrak{h}$, where each summand is $K$-invariant (not necessarily irreducible). Then, there exists a 2-parameter family of homogeneous metrics on $\mathsf{G}/\mathsf{K}$ with each parameter controls $\mathbb{S}^{d_{1}}$ and $\mathcal{Q}^{d_{2}}$. We further assume that the Ricci endomorphism of $\mathsf{G}/\mathsf{K}$ is diagonal for such a 2-parameter family. The background metric $b$ used in this paper is rescaled from $\tilde{b}$, so that the homogeneous metric
(2.1) $f_{1}^{2}\left.b\right|_{\mathfrak{h}/\mathfrak{k}}+f_{2}^{2}\left.b\right|_{\mathfrak{g}/\mathfrak{h}}$
has the Ricci endomorphism $\mathrm{diag}(r_{1}I_{d_{1}},r_{2}I_{d_{2}})$, where
(2.2) $\begin{split}r_{1}&=\frac{d_{1}-1}{f_{1}^{2}}+A\frac{f_{1}^{2}}{f_{2}^{4}},\quad r_{2}=\frac{d_{2}-1}{f_{2}^{2}}-2\frac{d_{1}}{d_{2}}A\frac{f_{1}^{2}}{f_{2}^{4}}.\end{split}$
With a principal orbit $\mathsf{G}/\mathsf{K}$ as above, each $f_{i}$ is a function defined on the $1$ dimensional orbit space $[0,t_{*}]$. From [EW00], the shape operator of $\mathsf{G}/\mathsf{K}$ in $M$ is $L:=\mathrm{diag}\left(\frac{\dot{f}_{1}}{f_{1}}I_{d_{1}},\frac{\dot{f}_{2}}{f_{2}}I_{d_{2}}\right)$. The cohomogeneity one Einstein equation is
(2.3) $\frac{\ddot{f_{i}}}{f_{i}}-\left(\frac{\dot{f_{i}}}{f_{i}}\right)^{2}$ $=-\mathrm{tr}(L)\frac{\dot{f_{i}}}{f_{i}}+r_{i}-\Lambda,\quad i=1,2;$
(2.4) $d_{1}\frac{\ddot{f_{1}}}{f_{1}}+d_{2}\frac{\ddot{f_{2}}}{f_{2}}$ $=-\Lambda.$
Alternatively, (2.4) can be replaced by
(2.5) $\mathrm{tr}(L^{2})-\mathrm{tr}^{2}(L)+d_{1}r_{1}+d_{2}r_{2}-(n-1)\Lambda=0.$
A solution $(f_{1},f_{2})$ defines a cohomogeneity one Einstein metric
(2.6) $dt^{2}+f_{1}^{2}(t)\left.b\right|_{\mathfrak{h}/\mathfrak{k}}+f_{2}^{2}(t)\left.b\right|_{\mathfrak{g}/\mathfrak{h}}$
on $M$ if the following initial and terminal conditions are satisfied.
(2.7) $\lim\limits_{t\to 0}(f_{1},\dot{f_{1}},f_{2},\dot{f_{2}})=(0,1,f,0),\quad\lim\limits_{t\to t_{*}}(f_{1},\dot{f_{1}},f_{2},\dot{f_{2}})=(0,-1,\bar{f},0),\quad f,\bar{f}>0.$
Follow [Chi24]. Let $d\eta=\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}dt$. Define
$X_{1}:=\frac{\frac{\dot{f_{1}}}{f_{1}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},\quad X_{2}:=\frac{\frac{\dot{f_{2}}}{f_{2}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},\quad Y:=\frac{\frac{1}{f_{1}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},\quad Z:=\frac{\frac{f_{1}}{f_{2}^{2}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}}.$
Define functions
(2.8) $\begin{split}&G:=d_{1}X_{1}^{2}+d_{2}X_{2}^{2},\quad H:=d_{1}X_{1}+d_{2}X_{2},\\ &R_{1}:=(d_{1}-1)Y^{2}+AZ^{2},\quad R_{2}:=(d_{2}-1)YZ-\frac{2d_{1}}{d_{2}}AZ^{2}.\end{split}$
Equations (2.3) are transformed to
(2.9) $\begin{split}X_{i}^{\prime}&=X_{i}H\left(G+\frac{1-H^{2}}{n}-1\right)+R_{i}-\frac{1-H^{2}}{n},\quad i=1,2;\\ Y^{\prime}&=Y\left(H\left(G+\frac{1-H^{2}}{n}\right)-X_{1}\right),\\ Z^{\prime}&=Z\left(H\left(G+\frac{1-H^{2}}{n}\right)+X_{1}-2X_{2}\right).\\ \end{split}$
The conservation law (2.5) is transformed to the following algebraic surface.
(2.10) $\mathcal{C}:\frac{d_{1}d_{2}}{n}(X_{1}-X_{2})^{2}+d_{1}R_{1}+d_{2}R_{2}=1-\frac{1}{n}.$
Equivalently, we have
(2.11) $\mathcal{C}:\frac{1}{n-1}\left(G-H^{2}+d_{1}R_{1}+d_{2}R_{2}\right)=\frac{1-H^{2}}{n}.$
By (2.9), we can set $Y$ and $Z$ to be non-negative without loss of generality. Furthermore, since
(2.12) $H^{\prime}=(H^{2}-1)\left(G+\frac{1}{n}(1-H^{2})\right)=(H^{2}-1)\left(\frac{1}{n}+\frac{d_{1}d_{2}}{n}(X_{1}-X_{2})^{2}\right),$
the set
$\mathcal{E}:=\mathcal{C}\cap\{Y,Z\geq 0\}\cap\{H^{2}\leq 1\}$
is flow invariant. We consider the dynamical system (2.9) on $\mathcal{E}$. Observe that the subsystem restricted on $\mathcal{E}\cap\{H=1\}$ is the cohomogeneity one Ricci-flat equation with the same group triple. To retrieve the original system, we apply [Chi24, Equation (2.10)].
3. Local Analysis
We perform local analysis at some important critical points of (2.9). The initial and terminal conditions (2.7) are respectively transformed to critical points
$p_{0}^{\pm}=\left(\pm\frac{1}{d_{1}},0,\frac{1}{d_{1}},0\right).$
The local existence of cohomogeneity one Einstein metrics is obtained from the local analysis at $p_{0}^{+}$. The linearization at $p_{0}^{+}$ is
$\begin{bmatrix}\frac{4d_{2}-2d_{1}(d_{2}-1)}{d_{1}n}&\frac{d_{2}(d_{1}-1)(d_{1}-d_{2})}{d_{1}^{2}n}&\frac{2(d_{1}-1)}{d_{1}}&0\\ \frac{2d_{1}}{n}&\frac{1}{d_{1}}+\frac{d_{2}-d_{1}}{n}&0&\frac{d_{2}-1}{d_{1}}\\ \frac{2d_{2}}{d_{1}n}&\frac{d_{2}(d_{2}-d_{1})}{d_{1}^{2}n}&0&0\\ 0&0&0&\frac{2}{d_{1}}\end{bmatrix}$
The critical point is hyperbolic. The following are the only two unstable eigenvectors that are tangent to $\mathcal{E}$. Both eigenvectors have the same eigenvalue as $\frac{2}{d_{1}}$.
$v_{1}=\begin{bmatrix}-2d_{2}(d_{2}-1)\\ 2d_{1}(d_{2}-1)\\ -d_{2}(d_{2}-1)\\ 2d_{1}(d_{1}+1)\end{bmatrix},\quad v_{2}=\begin{bmatrix}-d_{1}^{2}+d_{1}d_{2}-n\\ -2d_{1}^{2}\\ -d_{2}\\ 0\end{bmatrix}$
From the unstable version of [CL55, Theorem 4.5, Chapter 13], there exists a continuous $1$-parameter family $\gamma_{s}$ that emanates $p_{0}^{+}$ such that
(3.1) $\gamma_{s}\sim p_{0}^{+}+se^{\frac{2}{d_{1}}\eta}v_{1}+e^{\frac{2}{d_{1}}\eta}v_{2}+O\left(e^{\left(\frac{2}{d_{1}}+\delta\right)\eta}\right).$
The integral curve $\gamma_{0}$ is in the invariant set $\mathcal{E}\cap\{Z=0\}$. Each $\gamma_{s}$ with $s>0$ is in the interior :=$\mathcal{E}^{\circ}:=\mathcal{E}\cap\{Y,Z>0\}\cap\{H^{2}<1\}$ and it represents a locally defined positive Einstein metric around $\mathcal{Q}^{d_{2}}$. As $v_{1}$ is tangent to $\mathcal{RF}^{+}$, the integral curve that emanates from $p_{0}^{+}$ along $v_{1}$ represents a locally defined Ricci-flat metric around $\mathcal{Q}^{d_{2}}$.
To show the existence of a complete Einstein metric on $M$ is equivalent to showing a $\gamma_{s}$ joins $p_{0}^{\pm}$ for some $s>0$. The existence of such a heterocline depends on the existence of critical points with non-zero $Y$ and $Z$, which satisfy the quadratic equation
(3.2) $\frac{1}{Y^{2}}(R_{1}-R_{2})=\frac{n+d_{1}}{d_{2}}Al^{2}-(d_{2}-1)l+(d_{1}-1)=0,\quad l=\frac{Z}{Y}.$
Let $\Delta$ be the discriminant of (3.2). If $\Delta>0$, or equivalently,
(3.3) $A<\frac{1}{n+d_{1}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)},$
there are two real roots $\mu_{1}>\mu_{2}$ of the quadratic equation and hence two $\mathsf{G}$-invariant Einstein metrics on $\mathsf{G}/\mathsf{K}$. Subsequently, the following critical points exist.
$q_{i}^{\pm}=\left(\pm\frac{1}{n},\pm\frac{1}{n},y_{i},z_{i}\right),\quad R_{1}(y_{i},z_{i})=R_{2}(y_{i},z_{i})=\frac{n-1}{n^{2}},\quad i=1,2.$
Note that $\mu_{i}=\frac{z_{i}}{y_{i}}$. As $\mu_{1}$ and $\mu_{2}$ are roots for (3.2), we immediately obtain the following identities:
(3.4) $\mu_{1}+\mu_{2}=\frac{d_{2}(d_{2}-1)}{(n+d_{1})A},\quad\mu_{1}\mu_{2}=\frac{d_{2}(d_{1}-1)}{(n+d_{1})A}.$
Correspondingly, there exist two sine cone solutions of the cohomogeneity one Einstein equation that joins $q_{i}^{\pm}$, which are identified as the invariant sets
$\Phi_{i}:=\mathcal{E}\cap\{X_{1}-X_{2}=0\}\cap\{\mu_{i}Y-Z=0\},\quad i=1,2.$
A local analysis shows that the cone solution $\Phi_{2}$ is a focal attractor if
(3.5) $n\leq 8,\quad A<\frac{(9-n)(d_{2}n+7n+9d_{1})}{(d_{1}n-8n-9d_{1})^{2}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)}.$
For a $\mathsf{G}/\mathsf{K}$ that satisfies (3.5), one exploits the $\mathbb{Z}_{2}$-symmetry of the dynamical system to prove the existence of infinitely many cohomogeneity one Einstein metrics on $M$; see [Böh98, Corollary 5.8]. Unfortunately, there is no known example of a $\mathsf{G}/\mathsf{K}$ that satisfies (3.5) with $A>0$. We are hence motivated to find a new upper bound for $A$, presumably in the interval $\left(\frac{(9-n)(d_{2}n+7n+9d_{1})}{(d_{1}n-8n-9d_{1})^{2}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)},\frac{1}{n+d_{1}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)}\right)$, for the existence theorem.
Remark 3.1.
If $A\geq\frac{1}{n+d_{1}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)}$, there exists at most one $\mathsf{G}$-invariant Einstein metric on $\mathsf{G}/\mathsf{K}$. It is further deduced that the two summands type ansatz does not yield any cohomogeneity one Einstein metric on $M$ if such an inequality holds; see [Böh98, Theorem 3.1]. In an upcoming paper [Chi25a], we show that for some principal orbit, its characteristic triple satisfies (3.3) yet does not yield any Einstein metric on $M$ from the two summands type ansatz.
4. Global Analysis
In this section, we assume (3.3) so that the two real roots $\mu_{1}>\mu_{2}$ exist. Our global analysis looks for a new upper bound for $A$ to guarantee the existence of a $\gamma_{s}$ that joins $p_{0}^{\pm}$. From [Böh98, Theorem 4.6], the existence of such a heterocline is implied if a $\gamma_{s}$ has a large enough winding angle around $\Phi_{2}$.
Consider the following set whose boundary contains $p_{0}^{+}$ and $\Phi_{i}$.
+:=$\mathcal{E}^{+}:=\mathcal{E}\cap\{H\geq 0\}\cap\{X_{1}-X_{2}\geq 0\}\cap\{\mu_{1}Y-Z\geq 0\}.$
Proposition 4.1.
The set $\mathcal{E}^{+}$ is compact.
Proof.
By $\mu_{1}Y-Z\geq 0$, we have
(4.1) $\begin{split}d_{1}R_{1}+d_{2}R_{2}&=d_{1}(d_{1}-1)Y^{2}+d_{2}(d_{2}-1)YZ-d_{1}AZ^{2}\\ &\geq d_{1}(d_{1}-1)Y^{2}+\frac{d_{2}(d_{2}-1)-d_{1}A\mu_{1}}{\mu_{1}}Z^{2}.\end{split}$
By (3.4), we have $A=\frac{d_{2}(d_{2}-1)}{(n+d_{1})(\mu_{1}+\mu_{2})}\leq\frac{d_{2}(d_{2}-1)}{d_{1}\mu_{1}}$. The coefficient of $Z^{2}$ is positive. Since $H^{2}\leq 1$, each variable is bounded in $\mathcal{E}^{+}$ by (2.10). ∎
By (2.12), it is clear that an integral curve in $\mathcal{E}^{+}$ must escape the compact set through one of the three faces from the defining inequalities. Furthermore, by the derivatives of these defining functions, if an integral curve is not in $\mathcal{E}\cap\{Z=0\}$ or $\Phi_{i}$, it intersects $\partial\mathcal{E}^{+}$ transversely. Obviously, each $\gamma_{s}$ is initially in $\mathcal{E}^{+}$. Let $\sharp C(\gamma_{s})$ be the number of times that a $\gamma_{s}$ intersects
+$\mathcal{E}^{+}\cap\{H>0\}\cap\{X_{1}-X_{2}=0\}.$
We have the following proposition, which is an analogous statement to [Böh98, Theorem 4.6].
Lemma 4.2.
If there is an $s_{\bullet}>0$ such that $\sharp C(\gamma_{s_{\bullet}})\geq 1$, there exists some $s_{\star}\in(0,s_{\bullet})$ such that $\gamma_{s_{\star}}$ is a heterocline that joins $p_{0}^{\pm}$.
Proof.
Since $\gamma_{0}$ is in the invariant set $\mathcal{E}\cap\{Z=0\}$, we have
$(X_{1}-X_{2})^{\prime}=(X_{1}-X_{2})H\left(G+\frac{1-H^{2}}{n}-1\right)+(d_{1}-1)Y^{2}$
on $\gamma_{0}$. Hence, the function $X_{1}-X_{2}$ is positive along $\gamma_{0}$, and we have $C(\gamma_{0})=0$.
Each $\gamma_{s}$ with $s\in(0,s_{\bullet})$ escapes $\mathcal{E}^{+}$ transversely. Let $\gamma_{s}(\eta_{s})$ denote the escaping point, we have the following connected path
$\mathcal{P}:=\partial\mathcal{E}^{+}\cap\{\gamma_{s}(\eta_{s})\mid s\in[0,s_{\bullet}]\}$
that joins a point in +$\mathcal{E}^{+}\cap\{H=0\}\cap\{X_{1}-X_{2}>0\}$ (from $\gamma_{0}$) and a point in +$\mathcal{E}^{+}\cap\{H>0\}\cap\{X_{1}-X_{2}=0\}$ (from $\gamma_{s_{\bullet}}$). If $\mathcal{P}$ does not intersect +$\mathcal{E}^{+}\cap\{X_{1}=X_{2}=0\}$, the set necessarily intersects the integral curve $\Phi_{1}$, which is a contradiction. Therefore, there exists a $s_{\star}\in(0,s_{\bullet})$ for which $\gamma_{s_{\star}}$ escapes $\mathcal{E^{+}}$ through +$\mathcal{E}^{+}\cap\{X_{1}=X_{2}=0\}$. By the $\mathbb{Z}_{2}$-symmetry of (2.9), such a $\gamma_{s_{\star}}$ is a heterocline that joins $p_{0}^{\pm}$. ∎
In [Chi24], we prove the existence of $s_{\bullet}$ by constructing a compact barrier set $\mathcal{S}$ that forces each $\gamma_{s}$ in the set to intersect +$\mathcal{E}^{+}\cap\{H>0\}\cap\{X_{1}-X_{2}=0\}$. With a different structural triple $(d_{1},d_{2},A)$, the set $\mathcal{S}$ may or may not exist. The counterpart to $\mathcal{S}$ in this paper depends on a varying $(d_{1},d_{2},A)$. As $A$ falls below or exceeds certain thresholds determined by $(d_{1},d_{2})$, each $\gamma_{s}$ in the set is forced to escape through certain faces.
Define functions
(4.2) $\begin{split}S&:=\frac{Z}{Y}\frac{X_{1}+\frac{d_{2}}{2d_{1}}X_{2}}{X_{2}},\\ T&:=\frac{1-H^{2}}{n}\frac{1}{YZ},\\ P&:=X_{1}\left(R_{2}-\frac{1-H^{2}}{n}\right)-X_{2}\left(R_{1}-\frac{1-H^{2}}{n}\right)-2X_{2}(X_{1}-X_{2})\left(X_{1}+\frac{d_{2}}{2d_{1}}X_{2}\right).\end{split}$
The derivatives of functions (4.2) along $\gamma_{s}$ are
(4.3) $\begin{split}S^{\prime}&=-\frac{1}{X_{2}\left(X_{1}+\frac{d_{2}}{2d_{1}}X_{2}\right)}SP\\ T^{\prime}&=2TX_{2}\\ P^{\prime}&=P\left(H\left(3G+\frac{3}{n}(1-H^{2})-1\right)+\frac{n}{d_{1}}X_{2}-X_{1}\right)+(X_{1}-X_{2})(P+Q),\end{split}$
where
(4.4) $\begin{split}Q&=4X_{2}\left(X_{1}+\frac{d_{2}}{2d_{1}}X_{2}\right)\left(H+\frac{d_{2}}{2d_{1}}X_{2}\right)+\left(2X_{1}+2X_{2}+\frac{3d_{2}}{d_{1}}X_{2}\right)\frac{1-H^{2}}{n}\\ &\quad-2(d_{2}-1)X_{1}YZ-X_{2}\left(2(d_{1}-1)Y^{2}+\frac{3d_{2}}{d_{1}}(d_{2}-1)YZ\right).\end{split}$
Define
(4.5) $\begin{split}\mathfrak{S}_{A}(\sigma,\tau)&=\mathcal{E}^{+}\cap\{X_{1}\geq X_{2}\geq 0\}\cap\{P\geq 0\}\cap\{S\leq\sigma\}\cap\{T\geq\tau\},\end{split}$
where
$(\sigma,\tau)=\left(\frac{d_{2}(d_{2}-1)}{2d_{1}^{2}A},d_{2}-1-\frac{d_{1}+1}{d_{2}(d_{2}-1)}2d_{1}^{2}A\right).$
By Proposition 4.1, the set $\mathfrak{S}_{A}(\sigma,\tau)$ is compact.
Proposition 4.3.
If $A\in\left[0,\frac{d_{2}(d_{2}-1)^{2}}{2d_{1}^{2}(d_{1}+1)}\right)$, a non-trivial integral curve in $\mathfrak{S}_{A}(\sigma,\tau)$ either escapes the set through the face $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$, or $\mathfrak{S}_{A}(\sigma,\tau)\cap\{X_{1}=X_{2}>0\}$.
Proof.
Suppose a non-trivial integral curve escapes through the face $\{X_{2}=0\}$. If $X_{1}=0$ at the escaping point, there is nothing to prove. If $X_{1}>0$ at the escaping point, the $Z$-coordinate must vanish as the function $S$ is non-increasing in $\mathfrak{S}_{A}(k,l)$. With $X_{2}=Z=0$, the inequality $P\geq 0$ becomes $-X_{1}\frac{1-d_{1}^{2}X_{1}^{2}}{n}\geq 0$, meaning that the escaping point can only be $p_{0}^{+}$, a contradiction.
By (3.4), we have
(4.6) $\sigma=\frac{d_{2}(d_{2}-1)}{2d_{1}^{2}A}=\frac{n+d_{1}}{2d_{1}^{2}}\frac{\mu_{1}+\mu_{2}}{2}<\frac{n+d_{1}}{2d_{1}}\mu_{1}.$
The inequality $S\leq\sigma<\frac{n+d_{1}}{2d_{1}}\mu_{1}$ in $\mathfrak{S}_{A}(\sigma,\tau)$ implies
$\frac{Z}{Y}\frac{n+d_{1}}{2d_{1}}<\frac{Z}{Y}\frac{X_{1}+\frac{d_{2}}{2d_{1}}X_{2}}{X_{2}}=S<\frac{n+d_{1}}{2d_{1}}\mu_{1}.$
Therefore, an integral curve in $\mathfrak{S}_{A}(\sigma,\tau)$ does not escape through the face $\{Z-\mu_{1}Y=0\}$.
Since $S$ is non-increasing in $\mathfrak{S}_{A}(\sigma,\tau)$, an integral curve in $\mathfrak{S}_{A}(\sigma,\tau)$ does not escape the set through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{S=\sigma\}\cap\{P>0\}$. The inequality $\tau>0$ holds if $A\in\left[0,\frac{d_{2}(d_{2}-1)^{2}}{2d_{1}^{2}(d_{1}+1)}\right).$ Since $T$ is non-decreasing in $\mathfrak{S}_{A}(\sigma,\tau)$, an integral curve in $\mathfrak{S}_{A}(\sigma,\tau)$ does not escape the set through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{T=\tau\}$. ∎
Motivated by the derivative formula (4.3) and Proposition 4.3, we study the sign of $Q$ when the function is restricted on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$. The set $\mathfrak{S}_{A}(\sigma,\tau)$ is equivalently defined as
(4.7) $\begin{split}\mathfrak{S}_{A}(\sigma,\tau)&=\left(\bigcup\limits_{(k,l)\in[0,1]\times[0,\mu_{1}]}\mathcal{E}_{k,l}\right)\cap\{P\geq 0\}\cap\{S\leq\sigma\}\cap\{T\geq\tau\},\end{split}$
where
(4.8) $\begin{split}\mathcal{E}_{k,l}:=\mathcal{E}^{+}\cap\{X_{2}-kX_{1}=0,X_{1}\geq 0\}\cap\{Z-lY=0,Y\geq 0\}.\end{split}$
Parametrize the algebraic surface $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ using $(k,l)$ as in (4.8). The function $Q$ on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ is realized as a function on $\{(k,l)\mid(k,l)\in[0,1]\times[0,\mu_{1}]\}$. Specifically, with (2.11), the functions $P$ and $Q$ are equivalently defined as homogeneous polynomials of $(X_{1},X_{2},Y,Z)$ of degree $3$. On each slice $\mathcal{E}_{k,l}$, we have
(4.9) $\begin{split}P&=P_{Y}Y^{2}X_{1}+P_{X}X_{1}^{3},\quad Q=Q_{Y}Y^{2}X_{1}+Q_{X}X_{1}^{3},\end{split}$
where
(4.10) $\begin{split}P_{Y}(A,k,l)&=-\frac{d_{1}(n+d_{1}-2)+d_{2}(n+d_{1}-1)k}{d_{2}(n-1)}Al^{2}\\ &\quad+\frac{(d_{2}-1)(d_{1}+d_{2}k-1)}{n-1}l-\frac{(d_{1}-1)(d_{1}+d_{2}k-k)}{n-1},\\ P_{X}(k)&=\frac{d_{2}(d_{2}-1)(d_{1}-1)k^{2}-d_{1}d_{2}k^{2}+2(d_{2}-1)(d_{1}-1)d_{1}k+d_{1}^{2}(d_{1}-1)}{d_{1}(n-1)}(1-k),\\ Q_{Y}(A,k,l)&=-\frac{2nk+d_{2}k+2d_{1}}{n-1}Al^{2}\\ &\quad-\frac{2d_{1}(d_{1}-1)+d_{1}d_{2}k-3d_{2}k}{d_{1}(n-1)}(d_{2}-1)l+\frac{(2d_{1}+d_{2}k+2k)(d_{1}-1)}{n-1},\\ Q_{X}(k)&=4k\left(1+\frac{d_{2}k}{2d_{1}}\right)\left(d_{1}+d_{2}k+\frac{d_{2}k}{2d_{1}}\right)+\left(2+2k+\frac{3d_{2}k}{d_{1}}\right)\frac{d_{1}+d_{2}k^{2}-(d_{1}+d_{2}k)^{2}}{n-1}.\end{split}$
Proposition 4.4.
For any $d_{2}\geq d_{1}\geq 2$, the inequality $P_{X}\geq 0$ holds on $[0,1]$ and only vanishes at $k=1$.
Proof.
The quadratic factor of $P_{X}$ satisfies the inequality
(4.11) $\begin{split}&d_{2}(d_{2}-1)(d_{1}-1)k^{2}-d_{1}d_{2}k^{2}+2(d_{2}-1)(d_{1}-1)d_{1}k+d_{1}^{2}(d_{1}-1)\\ &\geq d_{2}(d_{2}-1)(d_{1}-1)k^{2}+(2(d_{2}-1)(d_{1}-1)d_{1}-d_{1}d_{2})k+d_{1}^{2}(d_{1}-1).\end{split}$
The coefficient of each $k^{i}$ is positive on $[0,1]$ since $d_{1},d_{2}\geq 2$. Therefore, the polynomial is positive on $[0,1]$. ∎
Define $\omega=Q_{Y}P_{X}-Q_{X}P_{Y}$. On $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$, we have
(4.12) $\begin{split}Q&=Q_{Y}Y^{2}X_{1}+Q_{X}\frac{P-P_{Y}Y^{2}X_{1}}{P_{X}}=\frac{\omega}{P_{X}}Y^{2}X_{1}.\end{split}$
The function $\omega$ has the same sign as $Q$ on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$. The specific formula for $\omega$ is
(4.13) $\begin{split}\omega(A,k,l)&=\frac{1}{d_{1}^{2}(n-1)}\left(\frac{2d_{1}+d_{2}k}{d_{2}}\omega_{2}Al^{2}+(d_{2}-1)k\omega_{1}l-(d_{1}-1)k^{2}\omega_{0}\right),\end{split}$
where
(4.14) $\begin{split}\omega_{2}(k)&=(2d_{1}^{2}d_{2}^{2}-d_{1}d_{2}^{3}+d_{2}^{3}-d_{2}^{2})k^{3}+(4d_{1}^{3}d_{2}-4d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}+4d_{1}d_{2}^{2}-2d_{1}d_{2})k^{2}\\ &\quad+(2d_{1}^{4}-5d_{1}^{3}d_{2}-2d_{1}^{3}+5d_{1}^{2}d_{2})k-2d_{1}^{4}+2d_{1}^{3},\\ \omega_{1}(k)&=(d_{1}d_{2}^{3}-4d_{1}d_{2}^{2}-d_{2}^{3}+3d_{2}^{2})k^{3}+(2d_{1}^{2}d_{2}^{2}-8d_{1}^{2}d_{2}+8d_{1}d_{2}-2d_{2}^{2})k^{2}\\ &\quad+(d_{1}^{3}d_{2}-4d_{1}^{3}+5d_{1}^{2}d_{2}+4d_{1}^{2}-6d_{1}d_{2})k+4d_{1}^{3}-4d_{1}^{2},\\ \omega_{0}(k)&=(d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-4d_{1}^{2}+4d_{1}d_{2})k+d_{1}^{3}d_{2}-d_{1}^{2}d_{2}+4d_{1}^{2}.\end{split}$
The critical point $p_{0}^{+}$, whose $(k,l)$-coordinate is $(0,0)$, is on the boundary of $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$. If $A=0$, the product Einstein metric is represented by the algebraic curve $\mathcal{E}\cap\{X_{2}=0\}\cap\left\{R_{2}-\frac{1-H^{2}}{n}=0\right\}$, at which both $P$ and $Q$ vanish. This is reflected by $\omega(0,0,l)$ being identically zero.
For any $A\geq 0$, both $\omega(A,k,l)$ and its gradient vanish at $(0,0)$. To learn the sign of $\omega(A,k,l)$ near $p_{0}^{+}$, we compute the Hessian of the function at $(0,0)$.
(4.15) $\left.\mathrm{Hess}(\omega)\right|_{(0,0)}=\begin{bmatrix}-\frac{2(d_{1}-1)(d_{1}d_{2}-d_{2}+4)}{n-1}&\frac{4(d_{1}-1)(d_{2}-1)}{n-1}\\ \frac{4(d_{1}-1)(d_{2}-1)}{n-1}&-\frac{8d_{1}^{2}(d_{1}-1)}{d_{2}(n-1)}A\end{bmatrix}.$
By the second partial test, we have the following proposition.
Proposition 4.5.
If $A\in\left[0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)$, the critical point $(0,0)$ is a saddle. If $A>\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$, the critical point $(0,0)$ is a local maximum.
Note that $\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$ is smaller than the upper bound in Proposition 4.3 if $d_{2}\geq d_{1}\geq 2$. By the derivative formula (4.3), Proposition 4.3 and Proposition 4.5 indicate that if any integral curve in $\mathfrak{S}_{A}(\sigma,\tau)$ does not escape the set only through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$, it is necessary that $A\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)$. We proceed to demonstrate that under this bound for $A$, there exists a $\gamma_{s}$ that is initially in $\mathfrak{S}_{A}(\sigma,\tau)$.
Proposition 4.6.
If $\omega$ is not identically zero along a $\gamma_{s}$ with $s>0$, the function $\omega$ has the same sign as $P$ once the integral curve leaves $p_{0}^{+}$.
Proof.
We have
(4.16) $\begin{split}Q+P=\frac{\omega}{P_{X}}Y^{2}X_{1}+\frac{P_{X}+Q_{X}}{P_{X}}P.\end{split}$
For any parameter $s\geq 0$, the integral curve $\gamma_{s}$ initially enters $\mathcal{E}^{+}$ and $P_{X}>0$ holds. From Proposition 5.1 we know that $P_{X}+Q_{X}<0$ for each $k\in[0,1]$. Assume $\omega>0$ once $\gamma_{s}$ is away from $p_{0}^{+}$, and suppose $P\leq 0$ initially. The inequality $Q+P>0$ holds initially by (4.16). As $d_{1}\geq 2$, we have
(4.17) $\begin{split}&\left(H\left(3G+\frac{3}{n}(1-H^{2})-1\right)+\frac{n}{d_{1}}X_{2}-X_{1}\right)(\gamma_{s})\\ &\sim\frac{2}{d_{1}}-1-((12+6d_{2})+4(d_{2}-1)(2d_{2}-1)s)e^{\frac{2}{d_{1}}\eta}+O\left(e^{\left(\frac{2}{d_{1}}+\delta\right)\eta}\right)\\ &<0\end{split}$
near $p_{0}^{+}$. Hence, the derivative $P^{\prime}$ in (4.3) is initially positive, which is a contradiction. The proof of the case $\omega<0$ is almost verbatim. ∎
Proposition 4.7.
Let $s_{\bullet}=\frac{d_{1}}{\tau}=\frac{d_{1}\sigma}{(d_{2}-1)\sigma-(d_{1}+1)}$. The integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(k,l)$ if $A\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)$.
Proof.
By (3.1), we have
$\lim\limits_{\eta\to-\infty}S(\gamma_{s_{\bullet}})=\sigma,\quad\lim\limits_{\eta\to-\infty}T(\gamma_{s_{\bullet}})=\tau.$
By the monotonicity of $S$ and $T$, the inequalities $S\leq\sigma$ and $T\geq\tau$ hold along $\gamma_{s_{\bullet}}$ once the integral curve leaves $p_{0}^{+}$. Since $\gamma_{s_{\bullet}}$ is initially tangent to $\{S=\sigma\}$, the $(k,l)$-projection of $\gamma_{s_{\bullet}}$ is tangent to the level curve $\omega(A,k,l)=0$ at $(0,0)$ along the vector $v=(1,\sigma)$. By (4.15), we have $\left.\mathrm{Hess}(\omega)\right|_{(0,0)}(v,v)>0$ given that $A<\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$. By Proposition 4.6, the inequality $P>0$ holds initially along $\gamma_{s_{\bullet}}$. Therefore, the integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(\sigma,\tau)$. ∎
Proposition 4.7 provides a condition on $A$ to form a barrier for $\gamma_{s_{\bullet}}$ locally, which means that $\gamma_{s_{\bullet}}$ may still escape $\mathfrak{S}_{A}(\sigma,\tau)$ through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ eventually. In the following, we explore the condition on $A$ to make $\mathfrak{S}_{A}(\sigma,\tau)$ a global barrier so that $\gamma_{s_{\bullet}}$ can only escape through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{X_{1}=X_{2}>0\}$.
Consider $\omega(A,k,l)$ as a quadratic function of $l$. We have the following proposition for the coefficients of $\omega$.
Proposition 4.8.
For any $d_{2}\geq d_{1}\geq 2$, the inequalities $\omega_{2}<0$ and $\omega_{0}>0$ hold for any $k\in[0,1]$.
Proof.
Since
$\omega_{2}\left(-\frac{d_{1}}{d_{2}}\right)=-\frac{d_{1}^{3}}{d_{2}}<0,\quad\frac{d\omega_{2}}{dk}\left(-\frac{d_{1}}{d_{2}}\right)=d_{1}^{2}(2d_{1}+1)>0,$
$\omega_{2}(0)=-2d_{1}^{3}(d_{1}-1)<\omega_{2}\left(-\frac{d_{1}}{d_{2}}\right),$
the function $\omega_{2}$ changes monotonicity at least once on $\left[-\frac{d_{1}}{d_{2}},0\right]$. Since $\omega_{2}$ changes monotonicity at most twice, the function decreases and then increases, or monotonically increases on $[0,1]$. From $\omega_{2}(1)=-d_{2}(n-1)(d_{1}n-n-d_{1})<0$ we know that $\omega_{2}<0$ on $[0,1]$.
For $\omega_{0}$, we have
$\frac{d\omega_{0}}{dk}(0)=2d_{1}d_{2}(d_{1}d_{2}-d_{1}-d_{3})+4d_{1}(d_{2}-d_{1})\geq 0.$
The function either monotonically increases, or increases and then decreases on $[0,1]$. Since
$\omega_{0}(0)=(d_{1}-1)d_{1}^{2}d_{2}+4d_{1}^{2}>0,\quad\omega_{0}(1)=d_{2}(n-1)(d_{1}n-n-d_{1})>0,$
the positivity of $\omega_{0}$ on $[0,1]$ is established. ∎
By Proposition 4.8, we learn that for each fixed $A>0$ and $k\in(0,1)$, the quadratic function $\omega(A,k,l)$ is concave down. By (4.12) and Proposition 4.4, the positivity of $Q$ on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}\cap\{X_{1}>X_{2}>0\}$ is implied by the following condition: for each $k\in(0,1)$, the quadratic $\omega(A,k,l)$ has two real roots $\mu_{2}(A,k)<\mu_{1}(A,k)$ and the interval for $l$ obtained from $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ is contained in $(\mu_{2}(A,k),\mu_{1}(A,k))$.
In the following, we specify how $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ gives rise to a closed interval for $l$. In terms of $k$ and $l$, the inequality $S\leq\sigma$ is equivalent to
(4.18) $l\leq\nu_{1}(A,k):=\sigma\frac{2d_{1}k}{2d_{1}+d_{2}k}=\frac{d_{2}(d_{2}-1)}{d_{1}A}\frac{k}{2d_{1}+d_{2}k},$
and it gives the upper bound for $l$. Presumably the lower bound $\nu_{2}(A,k)$ for $l$ is obtained from $\{P=0\}\cap\{T\geq\tau\}$. It takes the following computations to write $T\geq\tau$ in terms of $k$ and $l$. From (2.11), the inequality $T\geq\tau$ is equivalent to
$G-H^{2}+d_{1}R_{1}+d_{2}R_{2}-(n-1)\tau YZ\geq 0.$
Rewrite the above inequality as $T_{X}(k)X_{1}^{2}+T_{Y}(A,k,l)Y^{2}\geq 0,$ where $T_{X}$ and $T_{Y}$ are polynomials. With $P=0$, the condition $T\geq\tau$ is equivalent to
(4.19) $\begin{split}T_{X}X_{1}^{2}+T_{Y}Y^{2}=\frac{T_{Y}P_{X}-T_{X}P_{Y}}{P_{X}}Y^{2}\geq 0.\end{split}$
Define $\zeta(A,k,l)=T_{Y}P_{X}-T_{X}P_{Y}$. The specific formula for $\zeta(A,k,l)$ is
(4.20) $\zeta(A,k,l)=\zeta_{2}Al^{2}+\zeta_{1}l+\zeta_{0},$
where
(4.21) $\begin{split}\zeta_{2}(k)&=-\frac{(d_{1}+d_{2}k)(d_{1}+d_{2}k-1)(2d_{1}+d_{2}k)}{d_{2}}\\ \zeta_{1}(k,\tau)&=\frac{d_{2}-1}{d_{1}}(d_{1}+d_{2}k)(d_{2}k^{2}+d_{2}(d_{1}-1)k+d_{1}(d_{1}-1))\\ &\quad-\frac{1-k}{d_{1}}((d_{1}d_{2}^{2}-2d_{1}d_{2}-d_{2}^{2}+d_{2})k^{2}+(2d_{1}^{2}d_{2}-2d_{1}^{2}-2d_{1}d_{2}+2d_{1})k+d_{1}^{2}(d_{1}-1))\tau\\ \zeta_{0}(k)&=-k(d_{1}-1)(d_{1}+d_{2}k)(d_{1}+d_{2}k+1-2k).\end{split}$
As $P_{X}>0$ for each $k\in(0,1)$, the function $T-\tau$ and $\zeta$ have the same sign. We show in the following proposition that for any $A\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)$, the condition $\zeta(A,k,l)\geq 0$ always yields the lower bound for $l$ with each fixed $k\in(0,1)$.
Proposition 4.9.
For each $(A,k)\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)\times(0,1)$, the quadratic function $\zeta$ has two positive real roots. The smaller real root is in $(0,\nu_{1}(A,k))$.
Proof.
A straightforward computation shows
(4.22) $\begin{split}\zeta\left(A,k,\nu_{1}(A,k)\right)=\frac{k^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)A}(d_{2}(d_{2}-1)^{2}\beta_{0}-d_{1}^{2}\beta_{1}A),\end{split}$
where
(4.23) $\begin{split}\beta_{0}(k)&=(d_{1}d_{2}^{2}-2d_{1}d_{2}-d_{2}^{2}+d_{2})k^{2}+(2d_{1}^{2}d_{2}-2d_{1}^{2}-d_{1}d_{2}+2d_{1})k+d_{1}^{3}-d_{1},\\ \beta_{1}(k)&=(2d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+2d_{2})k^{2}\\ &\quad+(4d_{1}^{3}d_{2}+2d_{1}^{2}d_{2}^{2}-4d_{1}^{3}-2d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+4d_{1}d_{2}+d_{2}^{2}+4d_{1}-2d_{2})k\\ &\quad+2d_{1}^{4}+d_{1}^{3}d_{2}-2d_{1}^{2}d_{2}+2d_{1}^{2}+d_{1}d_{2}-4d_{1}.\end{split}$
The coefficients $\beta_{0}$ and $\beta_{1}$ are positive for each $k\in[0,1]$ by Proposition 5.2. Furthermore, from Proposition 5.3 we know that
$\min\limits_{k\in[0,1]}\left\{\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}}\frac{\beta_{0}}{\beta_{1}}\right\}=\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(n+d_{1})}\geq\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}.$
Therefore, the inequality $\zeta\left(A,k,\sigma\frac{2d_{1}k}{2d_{1}+d_{2}k}\right)>0$ is valid for any $(A,k)\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)\times(0,1)$. As $\zeta_{2},\zeta_{0}<0$ for any $k\in(0,1)$, we know that $\zeta(A,k,l)$ has two real roots for each $k\in(0,1)$. Furthermore, the smaller root is less than $\nu_{1}(A,k)$. ∎
Denote the smaller positive root of $\zeta(A,k,l)$ as $\nu_{2}(A,k)$. Proposition 4.9 guarantees the existence of the closed interval $[\nu_{2}(A,k),\nu_{1}(A,k)]$ for any $(A,k)\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)\times(0,1)$.
Proposition 4.10.
Define $A_{1}=\min\limits_{k\in[0,1]}\left\{\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)}\frac{\beta_{2}}{\omega_{0}}\right\}$, where
$\beta_{2}=(d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}+d_{2}^{2})k^{3}+(2d_{1}^{2}d_{2}^{2}-4d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}+2d_{1}d_{2})k^{2}+(d_{1}^{3}d_{2}-2d_{1}^{3}+d_{1}^{2}d_{2})k+2d_{1}^{3}.$
The number $A_{1}$ is positive and the function $\omega\left(A,k,\nu_{1}(A,k)\right)$ is positive for any $k\in(0,1)$ if $A\in(0,A_{1})$.
Proof.
With the help of Maple, we compute
(4.24) $\begin{split}&\omega\left(A,k,\nu_{1}(A,k)\right)=\frac{(d_{1}-1)k^{2}}{d_{1}^{2}(n-1)}\left(\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)}\frac{\beta_{2}}{A}-\omega_{0}\right)\\ \end{split}$
By Proposition 5.4 in the Appendix, the factor $\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)}\beta_{2}$ is positive for any $k\in[0,1]$. As $\omega_{0}$ is also positive by Proposition 4.8, we know that $A_{1}>0$. The proof is complete. ∎
Since $\omega(A,k,l)$ is concave down and $\omega(A,k,0)<0$ for any $k\in(0,1)$, Proposition 4.10 implies that an $\omega(A,k,l)$ with with $(A,k)\in(0,A_{1})\times(0,1)$ has two positive roots $\mu_{2}(A,k)<\mu_{1}(A,k)$ and $\nu_{1}(A,k)\in(\mu_{2}(A,k),\mu_{1}(A,k))$ for any $k\in(0,1)$.
With our choice of $(\sigma,\tau)$, it is clear that
$\nu_{2}(A,0)=\nu_{1}(A,0)=0,\quad\frac{d\nu_{2}}{dk}(A,0)=\frac{d\nu_{1}}{dk}(A,0)=\frac{d_{2}(d_{2}-1)}{2d_{1}^{2}A}.$
Therefore, the inclusion $[\nu_{2}(A,k),\nu_{1}(A,k)]\subset(\mu_{2}(A,k),\mu_{1}(A,k))$ holds with any sufficiently small $k>0$. Therefore, the inclusion can be extended to all $k\in(0,1)$ if $\nu_{2}(A,k)>\mu_{2}(A,k)$ holds for all $k\in(0,1)$. Such a condition puts another restriction on the upper bound for $A$.
Proposition 4.11.
There exists an $A_{2}>0$ such that for any $A\in(0,A_{2})$, the inequality $\nu_{2}(A,k)>\mu_{2}(A,k)$ holds for any $k\in(0,1)$.
Proof.
We compute the resultant of $(\omega,\zeta)$.
(4.25) $\begin{split}\mathrm{Res}_{l}(\omega,\zeta)&=-\frac{(d_{1}-1)(2d_{1}+d_{2}k)(1-k)^{2}k^{2}\beta_{3}^{2}}{d_{1}^{6}d_{2}^{3}(d_{2}-1)^{2}(n-1)^{2}}A\Theta_{d_{1},d_{2}},\end{split}$
where
(4.26) $\begin{split}\beta_{3}&=d_{2}(d_{1}-1)(d_{2}-2)k^{2}-d_{2}k^{2}+2d_{1}(d_{1}-1)(d_{2}-1)k+d_{1}^{2}(d_{1}-1)\\ &>-d_{2}k+2d_{1}(d_{1}-1)(d_{2}-1)k\\ &>0,\end{split}$
and
(4.27) $\begin{split}\Theta_{d_{1},d_{2}}(A,k)&=\theta_{2}A^{2}+\theta_{1}A+\theta_{0},\end{split}$
with coefficients $\theta_{i}(k)$ be polynomials; see Proposition 5.5. Furthermore, by Proposition 5.6, we have $\theta_{2}<0$ and $\theta_{0}\geq 0$ for any $k\in[0,1]$. The coefficient $\theta_{0}$ only vanishes at $k=0$. We also have
(4.28) $\begin{split}&\Theta_{d_{1},d_{2}}(A,0)\\ &=-(d_{1}^{3}d_{2}-d_{1}^{2}d_{2}+4d_{1}^{2})(4d_{1}^{4}-2d_{1}^{3})A^{2}+4d_{1}^{2}d_{2}(d_{2}-1)^{2}(d_{1}+1)(2d_{1}^{7}-2d_{1}^{5})A.\end{split}$
Therefore, for any fixed $k\in[0,1]$, the quadratic polynomial $\Theta_{d_{1},d_{2}}(A,k)$ always has exactly one positive root $A^{+}(k)$. Define $A_{2}=\min\limits_{k\in[0,1]}A^{+}(k)>0$. If $A\in\left(0,A_{2}\right)$, the factor $\Theta_{d_{1},d_{2}}(A,k)$ is positive for any $k\in[0,1]$. Therefore, the resultant $\mathrm{Res}_{l}(\omega,\zeta)$ does not vanish for any $(A,k)\in(0,A_{2})\times(0,1)$. As $\nu_{2}(A,k)>\mu_{2}(A,k)$ for a sufficiently small $k>0$, we conclude that $\nu_{2}(A,k)>\mu_{2}(A,k)$ for any $k\in(0,1)$. ∎
Define $\chi_{d_{1},d_{2}}=\min\left\{\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)},A_{1},A_{2}\right\}$. We obtain the following lemma.
Lemma 4.12.
Consider a structural triple $(d_{1},d_{2},A)$ with $A<\chi_{d_{1},d_{2}}$. The integral curve $\gamma_{s_{\bullet}}$ joins $p_{0}^{\pm}$ or $\sharp C(\gamma_{s_{\bullet}})\geq 1$.
Proof.
Since $\chi_{d_{1},d_{2}}\leq\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$, the integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(\sigma,\tau)$ by Proposition 4.7. Since $\chi_{d_{1},d_{2}}\leq A_{1},A_{2}$, the interval $[\nu_{2}(A,k),\nu_{1}(A,k)]$ exists and $[\nu_{2}(A,k),\nu_{1}(A,k)]\subset(\mu_{2}(A,k),\mu_{1}(A,k))$ for any $k\in(0,1)$ by Proposition 4.9, Proposition 4.10, and Proposition 4.11. Therefore, the inequality $Q>0$ holds on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}\cap\{X_{1}>X_{2}>0\}$. By (4.3) and Proposition 4.3, the integral curve $\gamma_{s_{\bullet}}$ intersects $\mathfrak{S}_{A}(\sigma,\tau)\cap\{X_{1}=X_{2}\geq 0\}$. ∎
Lemma 4.2 and Lemma 4.12 proves Theorem 1.1.
5. Appendix
Proposition 5.1.
The polynomial $P_{X}+Q_{X}$ is negative on $[0,1]$
Proof.
For $k\in[0,1]$ we have
(5.1) $\begin{split}-d_{1}^{2}(n-1)(P_{X}+Q_{X})&=(d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}-d_{2}^{3}+d_{1}d_{2}+d_{2}^{2})k^{3}\\ &\quad+(d_{1}^{2}d_{2}^{2}+2d_{1}^{2}d_{2}-d_{1}d_{2}^{2}+d_{1}d_{2})k^{2}\\ &\quad+(d_{1}^{3}d_{2}-d_{1}^{2}d_{2})k+(2d_{1}^{3}-2d_{1}^{2})(k-k^{2})+d_{1}^{3}(d_{1}-1)(1-k).\end{split}$
The last four terms above are non-negative and do not vanish simultaneously for any $k\in[0,1]$. If the coefficient of $k^{3}$ is non-negative, then the polynomial is positive on $[0,1]$. If the coefficient of $k^{3}$ is negative, we have
(5.2) $\begin{split}-d_{1}^{2}(n-1)(P_{X}+Q_{X})&>(d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}-d_{2}^{3}+d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(d_{1}^{2}d_{2}^{2}+2d_{1}^{2}d_{2}-d_{1}d_{2}^{2}+d_{1}d_{2})k^{2}\\ &=2d_{1}d_{2}(d_{1}d_{2}-d_{1}-d_{2})k^{2}+(d_{2}^{2}-d_{2})(d_{1}d_{2}-d_{1}-d_{2})k^{2}+d_{1}d_{2}k^{2}.\\ \end{split}$
Given $d_{2}\geq d_{1}\geq 2$, the polynomial above is non-negative on $[0,1]$. Hence, we have $P_{X}+Q_{X}<0$ on $[0,1]$. ∎
Proposition 5.2.
For any $d_{2}\geq d_{1}\geq 2$, the inequalities $\beta_{0}(k)>0$ and $\beta_{1}(k)>0$ hold for any $k\in[0,1]$.
Proof.
Since
(5.3) $\begin{split}\beta_{0}(k)&=(d_{1}d_{2}^{2}-2d_{1}d_{2}-d_{2}^{2}+d_{2})k^{2}+(2d_{1}^{2}d_{2}-2d_{1}^{2}-d_{1}d_{2}+2d_{1})k+d_{1}^{3}-d_{1},\\ &\geq(d_{1}-1)(d_{2}-1)d_{2}k^{2}-d_{1}d_{2}k+(2d_{1}^{2}d_{2}-2d_{1}^{2}-d_{1}d_{2}+2d_{1})k+d_{1}^{3}-d_{1}\\ &=(d_{1}-1)(d_{2}-1)d_{2}k^{2}+2d_{1}(d_{1}-1)(d_{2}-1)k+d_{1}^{3}-d_{1},\\ \end{split}$
we have $\beta_{0}>0$ on $[0,1]$.
Consider
(5.4) $\begin{split}\beta_{1}(k)&=(2d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+2d_{2})k^{2}\\ &\quad+(4d_{1}^{3}d_{2}+2d_{1}^{2}d_{2}^{2}-4d_{1}^{3}-2d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+4d_{1}d_{2}+d_{2}^{2}+4d_{1}-2d_{2})k\\ &\quad+2d_{1}^{4}+d_{1}^{3}d_{2}-2d_{1}^{2}d_{2}+2d_{1}^{2}+d_{1}d_{2}-4d_{1}.\\ \end{split}$
If $d_{1}=d_{2}=2$, then $\beta_{1}(k)=-12k^{2}+48k+36$ is positive on $[0,1]$. For $d_{2}\geq 3$, we argue that the coefficient of each $k^{i}$ is positive. The constant term of $\beta_{1}(k)$ is apparently positive. The coefficient of $k$ is
(5.5) $\begin{split}&4d_{1}^{3}d_{2}+2d_{1}^{2}d_{2}^{2}-4d_{1}^{3}-2d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+4d_{1}d_{2}+d_{2}^{2}+4d_{1}-2d_{2}\\ &=2d_{1}^{2}(2d_{1}d_{2}-2d_{1}-d_{2})+d_{1}d_{2}(2d_{1}d_{2}-3d_{2})+4d_{1}d_{2}+(d_{2}^{2}-2d_{2})+4d_{1},\end{split}$
which is positive for any $d_{2}\geq d_{1}\geq 2$. The coefficient of $k^{2}$ is
(5.6) $\begin{split}&2d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+2d_{2}\\ &=(2d_{2}^{2}-4d_{2})d_{1}^{2}+(d_{2}^{3}-2d_{2}^{2}-2d_{2})d_{1}-d_{2}^{3}+2d_{2}.\end{split}$
As a function of $d_{1}$, the expression above is positive at $d_{1}=2$ and increasing on $[2,\infty)$. Therefore, we have $\beta_{1}>0$ on $[0,1]$. ∎
Proposition 5.3.
$\min\limits_{k\in[0,1]}\left\{\frac{\beta_{0}}{\beta_{1}}\right\}=\frac{1}{n+d_{1}}.$
Proof.
We compute
(5.7) $\begin{split}\frac{d}{dk}\left(\frac{\beta_{0}}{\beta_{1}}\right)&=-\frac{(d_{1}-1)(d_{2}-2)}{\beta_{1}^{2}}\tilde{\beta},\end{split}$
where
(5.8) $\begin{split}\tilde{\beta}&=(2d_{1}^{2}d_{2}^{2}+2d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}+d_{2}^{2})k^{2}+(4d_{1}^{3}d_{2}+4d_{1}^{2}d_{2}^{2}-6d_{1}^{2}d_{2}-4d_{1}d_{2}^{2}+2d_{1}d_{2})k\\ &\quad+2d_{1}^{4}+2d_{1}^{3}d_{2}-4d_{1}^{3}-3d_{1}^{2}d_{2}+2d_{1}^{2}+d_{1}d_{2}.\end{split}$
The factor $\tilde{\beta}$ is positive as each coefficient for $k^{i}$ is positive given $d_{2}\geq d_{1}\geq 2$. Hence, we have $\min\limits_{k\in[0,1]}\left\{\frac{\beta_{0}}{\beta_{1}}\right\}=\frac{\beta_{0}}{\beta_{1}}(1)=\frac{1}{n+d_{1}}$. Note that if $(d_{1},d_{2})=(2,2)$, we have $\beta_{0}(k)=-2k^{2}+8k+6=\frac{1}{6}\beta_{1}(k)$. ∎
Proposition 5.4.
For any $d_{2}\geq d_{1}\geq 2$, we have $\beta_{2}(k)>0$ for any $k\in[0,1]$.
Proof.
We have
(5.9) $\begin{split}\beta_{2}(k)&=(d_{1}-1)(d_{2}-1)d_{2}^{2}k^{3}+d_{1}d_{2}^{2}(k^{2}-k^{3})+(2d_{1}^{2}d_{2}^{2}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+2d_{1}d_{2})k^{2}\\ &\quad+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k+2d_{1}^{3}(1-k)\\ &>(2d_{1}^{2}d_{2}^{2}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+2d_{1}d_{2})k^{2}+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k\\ &\geq(-4d_{1}^{2}d_{2}+2d_{1}d_{2})k^{2}+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k\\ &\geq(d_{1}^{3}d_{2}+2d_{1}d_{2}-3d_{1}^{2}d_{2})k\\ &\geq 0.\end{split}$
We conclude that $\beta_{2}>0$ on $[0,1]$. ∎
Proposition 5.5.
The coefficients for $\Theta_{d_{1},d_{2}}(A,k)$ is given by
(5.10) $\theta_{2}=4d_{1}^{4}(d_{1}+1)^{2}\theta_{2a}\theta_{2b},\quad\theta_{1}=\theta_{1a}\theta_{1b},\quad\theta_{0}=\theta_{0a}\theta_{0b},$
where
(5.11) $\begin{split}\theta_{2a}(k)&=(d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-4d_{1}^{2}+4d_{1}d_{2})k+d_{1}^{3}d_{2}-d_{1}^{2}d_{2}+4d_{1}^{2}\\ \theta_{2b}(k)&=(2d_{1}^{2}d_{2}^{2}-d_{1}d_{2}^{3}+d_{2}^{3}-d_{2}^{2})k^{3}\\ &\quad+(4d_{1}^{3}d_{2}-4d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}+4d_{1}d_{2}^{2}-2d_{1}d_{2})k^{2}\\ &\quad+(2d_{1}^{4}-5d_{1}^{3}d_{2}-2d_{1}^{3}+5d_{1}^{2}d_{2})k-2d_{1}^{4}+2d_{1}^{3}\\ \theta_{1a}(k)&=-4d_{1}^{2}d_{2}(d_{2}-1)^{2}(d_{1}+1)\\ \theta_{1b}(k)&=(d_{1}^{3}d_{2}^{5}-d_{1}^{2}d_{2}^{5}-4d_{1}^{3}d_{2}^{3}+3d_{1}^{2}d_{2}^{4}-d_{1}d_{2}^{4}+2d_{1}d_{2}^{3}-d_{2}^{4})k^{5}\\ &\quad+(4d_{1}^{4}d_{2}^{4}+4d_{1}^{4}d_{2}^{3}-6d_{1}^{3}d_{2}^{4}-16d_{1}^{4}d_{2}^{2}+18d_{1}^{3}d_{2}^{3}+2d_{1}^{2}d_{2}^{4}\\ &\quad\quad+4d_{1}^{3}d_{2}^{2}-10d_{1}^{2}d_{2}^{3}+8d_{1}^{2}d_{2}^{2}-6d_{1}d_{2}^{3})k^{4}\\ &\quad+(6d_{1}^{5}d_{2}^{3}+12d_{1}^{5}d_{2}^{2}-14d_{1}^{4}d_{2}^{3}-20d_{1}^{5}d_{2}+35d_{1}^{4}d_{2}^{2}+7d_{1}^{3}d_{2}^{3}\\ &\quad\quad+12d_{1}^{4}d_{2}-29d_{1}^{3}d_{2}^{2}+d_{1}^{2}d_{2}^{3}+8d_{1}^{3}d_{2}-12d_{1}^{2}d_{2}^{2})k^{3}\\ &\quad+(4d_{1}^{6}d_{2}^{2}+12d_{1}^{6}d_{2}-16d_{1}^{5}d_{2}^{2}-8d_{1}^{6}+28d_{1}^{5}d_{2}+8d_{1}^{4}d_{2}^{2}\\ &\quad\quad+8d_{1}^{5}-32d_{1}^{4}d_{2}+4d_{1}^{3}d_{2}^{2}-8d_{1}^{3}d_{2})k^{2}\\ &\quad+(d_{1}^{7}d_{2}+4d_{1}^{7}-9d_{1}^{6}d_{2}+8d_{1}^{6}+3d_{1}^{5}d_{2}-12d_{1}^{5}+5d_{1}^{4}d_{2})k\\ &\quad-2d_{1}^{7}+2d_{1}^{5}\\ \theta_{0a}(k)&=kd_{2}^{2}(d_{2}-1)^{4}(2d_{1}+d_{2}k)((2d_{1}d_{2}-2d_{1}+d_{2})k+2d_{1}^{2}+2d_{1})\\ \theta_{0b}(k)&=(2d_{1}^{2}-1)d_{2}^{2}k^{2}+(4d_{1}^{3}-2d_{1}^{2}-2d_{1})d_{2}k+2d_{1}^{3}(d_{1}-1).\end{split}$
Proof.
We use Maple to obtain the formula. ∎
Proposition 5.6.
For any $d_{2}\geq d_{1}\geq 2$, we have $\theta_{2}<0$ and $\theta_{0}\geq 0$ on $[0,1]$. The function $\theta_{0}$ only vanishes at $k=0$.
Proof.
From observation it is clear that $\theta_{0}=\theta_{0a}\theta_{0b}\geq 0$ on $[0,1]$ and only vanishes at $k=0$. For $\theta_{2}$ we have
(5.12) $\begin{split}\theta_{2a}&=d_{2}^{2}(d_{1}d_{2}-d_{1}-d_{2})k^{2}+(d_{2}^{2}-d_{1}d_{2})k^{2}+(d_{1}d_{2}^{2}+d_{1}d_{2})(k-k^{2})+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2})k\\ &\quad+(2d_{1}^{2}d_{2}-d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+3d_{1}d_{2}-3d_{1}^{2})k+(d_{1}^{2}d_{2}+d_{1}^{2})(1-k)+(d_{1}^{3}d_{2}-2d_{1}^{2}d_{2}+3d_{1}^{2}).\end{split}$
Given $d_{2}\geq d_{1}\geq 2$, each term above is non-negative on $[0,1]$. In particular, the last term is positive. Hence $\theta_{2a}>0$ on $[0,1]$.
For the other factor $\theta_{2b}$, we have
(5.13) $\begin{split}\theta_{2b}&=-(d_{1}d_{2}^{3}-d_{2}^{3}+d_{2}^{2})k^{3}-2d_{1}^{2}d_{2}^{2}(k^{2}-k^{3})\\ &\quad-(d_{1}^{3}d_{2}-3d_{1}^{2}d_{2}+2d_{1}d_{2})k^{2}-(2d_{1}^{2}d_{2}^{2}-4d_{1}d_{2}^{2})k^{2}\\ &\quad-(5d_{1}^{3}d_{2}-5d_{1}^{2}d_{2})(k-k^{2})-(2d_{1}^{4}-2d_{1}^{3})(1-k)\\ \end{split}.$
Given $d_{2}\geq d_{1}\geq 2$, all terms above are non-positive and do not vanish simultaneously for any $k\in[0,1]$. We have $\theta_{2b}<0$ on $[0,1]$. The proof is complete. ∎
Acknowledgement. The author would like to thank the conference organizers of “Einstein Spaces and Special Geometry” at the Institut Mittag-Leffler in July 2023, where part of this work was presented. The author thanks Christoph Böhm, Andrew Dancer, Lorenzo Foscolo, and Claude LeBrun for insightful discussions during the conference. Special thanks to Christoph Böhm for his advice in proving the inhomogeneity. The author also thanks McKenzie Wang and Xiping Zhu for valuable comments on an early draft of this paper.